# Near-Field SWIPT with gMIMO in the Upper Mid-Band: Opportunities, Challenges, and the Way Forward

### v1

@@ Abstract
Index Terms:
gigantic MIMO, near-field, simultaneous wireless information and power transfer, upper mid-band frequencies.
@@ Introduction
While SWIPT and MIMO have been extensively studied individually, their combination in the emerging near-field regime, particularly within the upper mid-band, remains an underexplored frontier. Most SWIPT studies focus on far-field scenarios where the plane-wave approximation simplifies analysis and implementation. These approaches often neglect the unique challenges and opportunities posed by near-field propagation, such as the need for spherical wave-based channel models and the precise alignment required for energy beamforming. Similarly, advancements in near-field MIMO systems have emphasized adaptive beamforming techniques and spatial resolution improvements but have yet to fully consider the implications of incorporating SWIPT into their design. This paper aims to address these gaps by providing a comprehensive and focused exploration of near-field SWIPT with gMIMO operating in the upper mid-band. To ensure widespread deployment, addressing scalability challenges and diverse environmental conditions is essential. This includes applications across rural, urban, and industrial settings.
@@
Figure 1: Near-field SWIPT use cases with gMIMO. On the top half, we present several IoE applications where SWIPT can play an essential role. On the bottom half: Selected use case for smart factory with near-field SWIPT gMIMO.
@@
Similar to conventional communication systems, SWIPT has largely been studied in the far-field [7]. Most current research on radio-frequency (RF) WPT focuses on systems operating in the sub-6 GHz band due to better propagation conditions and “large” isotropic receivers and utilizing a relatively small number of antennas. In practical scenarios (e.g., distances $>$ 11112222 meters), charging devices are generally positioned in the far-field region [8, 9]. However, future 6G-based devices are expected to be located in the radiative near-field of larger arrays [9] thanks to larger antenna aperture and higher operating frequency. In this context, SWIPT can also leverage beamfocusing and massive spatial multiplexing. When EH users and ID users are positioned separately, near-field beamfocusing enables the system to direct signals more precisely, serving ID users with less power leakage and allowing surplus energy to be more efficiently delivered to EH users. While EH users may also benefit from harvesting ambient RF energy from information-bearing beams, the ability to spatially separate users in the near field reduces interference and allows for better beam management. In another SWIPT configuration where users perform simultaneous EH and ID, beamfocusing facilitates optimized time- or power-splitting strategies, thereby maximizing the harvested energy under joint EH-ID constraints.
Applications
@@
SWIPT has been applied across a diverse range of domains, including the IoE domain [10]. The IoE seeks to interconnect a vast array of devices and systems through the integration of cloud/edge computing and big data, enabling advanced 6G use cases such as autonomous vehicles, drone swarms, smart cities, and smart agriculture. Although industrialized environments are the most common and popular use cases, broader applications, such as emergency communications and wearable healthcare devices, are particularly noteworthy, as shown on the top of Fig. 1. For instance, battery-less devices that harvest energy from broadcast signals during disasters can maintain connectivity in critical times.
@@
The emergence of gMIMO in the upper mid-band presents a novel approach by accommodating a larger number of antennas within a given form factor, while avoiding the severe propagation challenges of millimeter-wave (mmWave) bands. In particular, the extended Fraunhofer distance—reaching up to tens of meters for gMIMO arrays in this band—enables effective near-field SWIPT in a wide range of scenarios, including indoor and predominantly outdoor applications. For instance, sensors deployed in factories, agricultural fields, and smart cities can leverage near-field SWIPT efficiently, as illustrated in Fig. 1.
In smart cities featuring autonomous vehicles and Internet of Vehicles (IoV) networks, roadside units equipped with gMIMO arrays can significantly enhance data rates while simultaneously delivering power to recharge sensor nodes within the IoV network. Similarly, in smart factories with numerous EH users, gMIMO presents an unparalleled opportunity to serve substantially more users within the same time-frequency resources. After addressing the communication demands of multiple users, gMIMO can still support a large number of EH users, fostering a sustainable IoE ecosystem. Smart farming and agricultural sites would be one of the ideal applications where sensor nodes in agricultural fields can act as EH users while “agribots” serve as information decoders.
@@
Fig. 2: $16$$\times$$4$, $40$$\times$$10$, $80$$\times$$20$
Figs. 3 and 4: $40$$\times$$10$, $80$$\times$$20$
Fig. 5: $40$$\times$$10$
@@
Noise power spectral density $-204$ dBW/Hz
Noise figure $7$ dB
Uplink pilot power $100$ $\mu$W
Number of symbols in a coherence block 10 0001000010\,00010 000
Analog circuit power of EH devices $3$ mW
@@
The initial 5G deployments operate at a frequency of $3.5$ GHz. However, frequencies in the upper mid-band, such as $6.425$-$7.125$ GHz, have also been identified as potential candidates for next-generation wireless systems. To meet the ever-growing data rate demands of applications like immersive reality and holographic communications, researchers are exploring even higher frequencies [6]. mmWaves spanning $24$ GHz and above are considered a natural extension due to the abundance of available spectrum in these ranges. Consequently, 5G research primarily focused on two frequency ranges: FR1 ($410$-$7125$ MHz) and FR2 (24.2524.2524.2524.2571.071.071.071.0 GHz) [11].
@@
While FR1 is already heavily utilized and crowded, FR2 faces significant practical challenges, including high susceptibility to blockages and inefficient RF hardware. An alternative is emerging in the form of the upper mid-band frequencies (7.1257.1257.1257.12524242424 GHz), also known as FR3. The upper mid-band offers an attractive compromise, providing more bandwidth than FR1 and better propagation characteristics than FR2. Moreover, this range supports a greater number of antenna elements without altering the physical size of the whole array, still resulting in near-field effect advantages for spatial multiplexing and beamfocusing. Additionally, employing multiple antennas at the receiver side can enhance the spatial degrees-of-freedom [4].
@@
Upper mid-band frequencies can meet both the performance goals of 6G and the unique demands of SWIPT systems by balancing energy harvesting and communication. Spatial multiplexing improves SE for ID users, while wireless power transfer primarily benefits from high beamforming gain and near-field focusing, which enable precise and efficient energy delivery to EH users. To this end, we analyze the EH capability of a SWIPT system under various configurations, including different frequency bands, numbers of ID users, and antenna configurations at the serving base station (BS) in Fig. 2. The path loss model for an indoor factory environment from [12] is used, excluding shadow fading. Correlated Rician fading with a $\kappa$-factor of $20$ dB and isotropic scattering is employed. 10 EH users are randomly deployed, while the horizontal axis in Fig. 2 denotes the number of ID users served simultaneously. The distances, azimuth angles, and elevation angles of all EH and ID users are randomly generated across multiple realizations. The distances and angles are sampled uniformly within the range $[1.5-25]$ meters and $[-\pi/2,\pi/2]$, respectively. The BS has a total power budget of $10$ W, with antennas arranged in a uniform planar array as indicated in the legend. The inter-antenna spacing is set to half a wavelength. The system operates over a bandwidth of $100$ MHz, and zero-forcing precoding is applied with perfect channel state information (CSI). A power allocation scheme ensures a minimum SE of $4$ bit/s/Hz for each ID user, with the remaining power budget used to transmit energy beams. These energy beams are designed to lie in the null space of the ID user channels to avoid interference. Although the current design employs separate beams for data and energy transmission, future research could investigate the joint optimization of information-carrying signals to simultaneously support EH, thus reducing the reliance on dedicated energy beams.
@@
Fig. 2 illustrates four configurations, with the first three maintaining the same physical array area but operating at increasing frequencies, thereby accommodating a higher number of antennas. The upper mid-band frequencies with a gMIMO array show significantly improved harvested power per EH user, particularly as the number of ID users increases. This is evident in the energy saved through more focused data transmission to ID users, allowing that energy to be used to serve more EH users. This gain stems from the fact that higher frequencies allow packing more antennas into the same physical array area, which results in narrower beams and higher array gain. Consequently, achieving the SE targets for ID users requires less transmit power, allowing a greater portion of the power budget to be allocated to energy transmission. In contrast, for mmWave bands (i.e., the case when $f_{c}=30$ MHz in Fig. 2), the number of antennas is capped at $80\times 20=1600$ due to hardware limitations. The reduced form factor at these higher frequencies restricts the advantages of adding more antennas, which diminishes the efficiency gains compared to the upper mid-band. While larger mmWave arrays could in theory achieve even higher gains, such arrays are less feasible in practice due to increased cost and complexity. Hence, gMIMO in the upper mid-band emerges as a promising solution, offering an optimal balance between hardware complexity and enhanced spatial multiplexing capabilities.
@@
Figure 2: Received power versus the number of ID users for different bands.
@@
To illustrate how near-field beamfocusing enhances channel properties and increases EH capability, Fig. 3 compares the far-field and near-field scenarios. For a fair comparison of the channel structure, all channel vectors are scaled to correspond to a path loss at a distance of $25$ m. Thus, the observed differences in Fig. 3 arise solely from the planar and spherical wave characteristics inherent to the far-field and near-field regions, respectively. All users are positioned within a narrow region near the broadside direction of the gMIMO array to highlight the advantages of near-field massive spatial multiplexing. As shown in Fig. 3, under identical path loss conditions, near-field propagation enables greater average harvested energy. This improvement stems from the more efficient utilization of transmit power and the increased spatial multiplexing capability offered by near-field focusing. With more focused beams, the system can more easily satisfy the SE requirements of ID users using less power, which leaves more power available for EH.
@@
For systems using the power splitting protocol, it is crucial to develop adaptive algorithms that dynamically adjust the power splitting ratio in real time based on battery levels, data requirements, and environmental factors. These algorithms can leverage machine learning to optimize the trade-off between power and data transfer continuously.
@@
The hardware costs of a gMIMO SWIPT system can be analyzed from both the BS and user terminal perspectives. At the BS, the presence of a gigantic number of antennas increases hardware complexity and processing costs compared to conventional massive MIMO systems. To fully exploit the potential of a gMIMO array, digital beamforming is essential, introducing additional challenges to hardware design and cost. On the user side, ID user terminals equipped with a greater number of antennas also experience increased hardware costs. Similarly, EH users can benefit from multiple rectennas—antennas that rectify signals— which enhance harvested energy but bring additional hardware costs.
@@
The upper mid-band frequencies exhibit sparser propagation characteristics compared to sub-6 GHz systems, which simplify channel estimation using parametric techniques. These methods estimate a smaller set of channel parameters rather than the entire channel, enabling reduced pilot utilization. For channels with a strong line of sight (LOS) component, even though more parameters are required compared to far-field channels, parametric estimation techniques still necessitate significantly fewer pilot resources. Classical direction-of-arrival estimation algorithms and compressed sensing-based methods can be effectively employed for this purpose.
@@
To observe the effect of channel estimation errors on the harvested energy, we consider a gMIMO system with $20$ ID users and $5$ EH users in Fig. 4. Under the requirement for SE of $4$ bit/s/Hz for the ID users, the cumulative distribution function (CDF) of the net average harvested power per EH user is plotted. Initially, the EH users transmit orthogonal pilot sequences of $5$ symbols to the gMIMO BS, which then estimates their channels using either a parametric maximum likelihood (ML)-based channel estimator or a least-squares (LS) channel estimator. The analog circuit power consumption of each EH device is set to $3$ mW, consistent with the low-power IoT device specifications in [15]. As shown in Fig. 4, the CDF curves corresponding to the ML estimator are shifted further to the right, indicating that more power can be harvested when using the ML estimator, as it provides higher channel estimation accuracy. However, this improvement comes at a significant computational cost due to the grid search required in ML estimation. Striking a balance between performance and computational complexity presents an interesting and important research direction.
@@
In the lower part of Fig. 5, we consider near-field SWIPT with a higher number of ID user antennas, as indicated in the legend. In this case, the location of the robotic arm is also randomized while maintaining a fixed distance of $25$ m. Interestingly, increasing the number of ID user antennas decreases the maximum average harvested power. This can be attributed to the fact that the energy beamformers are designed in the null space of the ID precoding matrix, which has a higher dimension as the number of subchannels increases in a near-field scenario. Consequently, the dimensionality of the space available for the energy precoder becomes more restricted when the number of antennas at the ID user increases. This observation is noteworthy, as the optimal number of antennas may vary depending on the SE requirement of the robotic arm. For lower SE requirements, fewer antennas can be activated, which leads to a modified version of the classical water-filling algorithm. This adaptation makes optimization for multi-antenna users an intriguing research direction.
@@
Sparse arrays have been extensively studied in the field of array signal processing. In a sparse array, sensors or antennas are arranged in a (non)uniform manner, with part of/all elements spaced more than half a wavelength apart. This design allows for an increased array aperture with a fixed number of antennas, improving the array resolution. Sparse arrays also offer a potential solution to the hardware cost challenges of gMIMO systems, where the goal is to deploy antennas within a given array aperture to preserve the near-field effects for SWIPT capabilities of the gMIMO array. By adopting sparse array designs, the cost can be reduced compared to dense antenna placement within the array aperture. Furthermore, sparse arrays can be dynamically configured by selecting antennas from a dense antenna grid based on real-time CSI and the locations of EH and ID users.
As previously mentioned, modular arrays (also referred to as modular MIMO) use groups of densely packed antennas arranged sparsely across the aperture. This structure is particularly suitable for environments like tall buildings. Beyond hardware cost reduction, modular arrays enable the development of novel distributed SWIPT precoding algorithms by leveraging cost-efficient local processing at each subarray. Future research should also explore advanced rectennas with higher RF-to-DC conversion efficiency to further reduce energy consumption while maintaining performance.

### v2

@@ Introduction
While SWIPT and MIMO have been extensively studied individually, their combination in the emerging near-field regime, particularly within the upper mid-band, remains an underexplored frontier. Most SWIPT studies focus on far-field scenarios where the plane-wave approximation simplifies analysis and implementation. These approaches often neglect the unique challenges and opportunities posed by near-field propagation. In parallel, research on near-field MIMO has primarily concentrated on enhancing spatial resolution and developing adaptive beamforming techniques, without fully examining the implications of incorporating SWIPT into these systems. This paper addresses these gaps by providing a comprehensive study of near-field SWIPT with gMIMO operating in the upper mid-band. It highlights how near-field beamfocusing and massive spatial multiplexing can be exploited to efficiently serve both information-decoding and energy-harvesting users. Furthermore, it analyzes precoding and power allocation strategies under realistic conditions, including shadowing, imperfect CSI, and non-linear energy harvesting effects. To bridge theory and practice, the study introduces detailed application scenarios—such as smart cities, factories, and agriculture—and presents a smart factory case study demonstrating the spectral efficiency and energy harvesting gains achievable through near-field beamfocusing.
@@
Figure 1: Near-field SWIPT with gMIMO in a selected use case for smart factory.
@@
Similar to conventional communication systems, SWIPT has largely been studied in the far-field [7]. Most current research on radio-frequency (RF) WPT focuses on systems operating in the sub-6 GHz band due to better propagation conditions and “large” isotropic receivers and utilizing a relatively small number of antennas. In practical scenarios (e.g., distances >> 1122 meters), charging devices are generally positioned in the far-field region [8, 9]. However, future 6G-based devices are expected to be located in the radiative near-field of larger arrays [9] thanks to larger antenna aperture and higher operating frequency. In this context, SWIPT can also leverage beamfocusing and massive spatial multiplexing. When EH users and ID users are positioned separately, near-field beamfocusing enables the system to direct signals more precisely, serving ID users with less power leakage and allowing surplus energy to be more efficiently delivered to EH users. While EH users may also benefit from harvesting ambient RF energy from information-bearing beams, the ability to spatially separate users in the near field reduces interference and allows for better beam management. In another SWIPT configuration where users perform simultaneous EH and ID, beamfocusing facilitates optimized time- or power-splitting strategies, thereby maximizing the harvested energy under joint EH-ID constraints.
Near-Field SWIPT Applications and Use Cases
@@
SWIPT has been applied across diverse domains, particularly within the IoE [10], which integrates cloud/edge computing and big data to enable advanced 6G use cases such as autonomous vehicles, drone swarms, smart cities, and smart agriculture. Beyond industrial applications, scenarios like emergency communications and wearable healthcare are also important, where battery-less devices can harvest energy from broadcast signals to maintain connectivity during disasters.
@@
The emergence of gMIMO in the upper mid-band offers a promising approach for enabling near-field SWIPT. By accommodating a large number of antennas within a given form factor, gMIMO avoids the severe propagation challenges of millimeter-wave (mmWave) bands while benefiting from extended Fraunhofer distances that can reach tens of meters. This property allows effective near-field operation across a wide range of indoor and outdoor scenarios. For example, sensors deployed in factories, agricultural fields, and smart cities can efficiently harvest energy while maintaining reliable communication links.
Smart cities and IoV networks: In urban environments with autonomous vehicles and Internet of Vehicles (IoV) networks, roadside gMIMO units can simultaneously deliver high data rates and recharge sensor nodes within the network. Use cases such as traffic management, smart street lighting control, and smart public transportation can greatly benefit from the spatial precision and simultaneous energy–information transfer enabled by near-field SWIPT. Moreover, unlike far-field line-of-sight (LOS) channels, near-field propagation enables spatial multiplexing even for users at similar angles but different distances. This capability allows gMIMO arrays to support multiple vehicular or infrastructure nodes simultaneously, enhancing spectral efficiency in dense urban deployments or making more power available for EH.
Smart factories: In industrial environments with numerous EH users, gMIMO enables the simultaneous support of many devices within the same time–frequency resources. After meeting the communication requirements of robotic systems and controllers, surplus power can be directed to battery-less sensors, fostering a sustainable IoE ecosystem.
Smart agriculture: In farming scenarios, near-field SWIPT can power sensors while maintaining reliable control links with mobile “agribots.” Typical use cases include machinery tracking and real-time quality monitoring using distributed sensors across large fields. Moreover, since the system operates in the upper mid-band, a large number of antenna elements can be integrated into a compact gMIMO array, allowing it to be deployed at the edge of fields without occupying excessive space. This compact yet powerful array can exploit spatial diversity in both distance and angle, enhancing link reliability and enabling simultaneous support for multiple distributed devices in large-scale agricultural deployments.
@@
[…]
@@
Noise power spectral density 204-204 dBW/Hz
Noise figure 77 dB
[…]
Number of symbols in a coherence block 10 00010\,000
Analog circuit power of EH devices 33 mW
@@
The initial 5G deployments operate at a frequency of $3.5$ GHz. However, frequencies in the upper mid-band, such as 6.4256.425-$7.125$ GHz, have also been identified as potential candidates for next-generation wireless systems. To meet the ever-growing data rate demands of applications like immersive reality and holographic communications, researchers are exploring even higher frequencies [6]. mmWaves spanning $24$ GHz and above are considered a natural extension due to the abundance of available spectrum in these ranges. Consequently, 5G research primarily focused on two frequency ranges: FR1 (410410-$7125$ MHz) and FR2 (24.2524.2571.071.0 GHz) [11].
@@
While FR1 is already heavily utilized and crowded, FR2 faces significant practical challenges, including high susceptibility to blockages and inefficient RF hardware. An alternative is emerging in the form of the upper mid-band frequencies (7.1257.1252424 GHz), also known as FR3. The upper mid-band offers an attractive compromise, providing more bandwidth than FR1 and better propagation characteristics than FR2. Moreover, this range supports a greater number of antenna elements without altering the physical size of the whole array, still resulting in near-field effect advantages for spatial multiplexing and beamfocusing. Additionally, employing multiple antennas at the receiver side can enhance the spatial degrees-of-freedom [4].
@@
Upper mid-band frequencies can meet both the performance goals of 6G and the unique demands of SWIPT systems by balancing energy harvesting and communication. Spatial multiplexing improves SE for ID users, while wireless power transfer primarily benefits from high beamforming gain and near-field focusing, which enable precise and efficient energy delivery to EH users. To this end, we analyze the EH capability of a SWIPT system under various configurations, including different frequency bands, numbers of ID users, and antenna configurations at the serving base station (BS) in Fig. 2. The path loss model for an indoor factory environment from [12] is adopted, incorporating lognormal shadowing with a standard deviation of $4\,\text{ dB}$. The gMIMO array is strategically positioned on the ceiling to ensure an LOS connection to each user. Correlated Rician fading with isotropic scattering is considered, where the $\kappa$-factor follows a lognormal distribution with a mean of $7\,\text{ dB}$ and a standard deviation of $8\,\text{ dB}$. 10 EH users are randomly deployed, while the horizontal axis in Fig. 2 denotes the number of ID users served simultaneously. The distances, azimuth angles, and elevation angles of all EH and ID users are randomly generated across multiple realizations. The distances and angles are sampled uniformly within the range [1.525][1.5-25] meters and $\pi/2,\pi/2]$, respectively. The BS has a total power budget of $10$ W, with antennas arranged in a uniform planar array as indicated in the legend. The inter-antenna spacing is set to half a wavelength. The system operates over a bandwidth of $100\,\mathrm{MHz}$, and zero-forcing precoding is employed under imperfect channel state information (CSI), where the least-squares (LS) estimator is used.
The performance is evaluated in terms of spectral efficiency (SE) for ID users and average harvested power for EH users. We examine different array configurations, operating frequencies (as summarized in Table I), and channel models (near-field vs. far-field) to assess their impact on energy transfer and spatial multiplexing capabilities. The specific parameter settings corresponding to each figure are listed in Table I.
A power allocation scheme ensures a minimum SE of $4$ bit/s/Hz for each ID user, with the remaining power budget used to transmit energy beams. These energy beams are designed to lie in the null space of the ID user channel estimates to reduce interference. Although the current design employs separate beams for data and energy transmission, future research could investigate the joint optimization of information-carrying signals to simultaneously support EH, thus reducing the reliance on dedicated energy beams.
@@
Fig. 2 illustrates four configurations, with the first three maintaining the same physical array area but operating at increasing frequencies, thereby accommodating a higher number of antennas. The upper mid-band frequencies with a gMIMO array show significantly improved harvested power per EH user, particularly as the number of ID users increases. This is evident in the energy saved through more focused data transmission to ID users, allowing that energy to be used to serve more EH users. This gain stems from the fact that higher frequencies allow packing more antennas into the same physical array area, which results in narrower beams and higher array gain. Consequently, achieving the SE targets for ID users requires less transmit power, allowing a greater portion of the power budget to be allocated to energy transmission. In contrast, for mmWave bands (i.e., the case when $f_{c}=30$ GHz in Fig. 2), the number of antennas is capped at $80\times 20=1600$ due to hardware limitations. The reduced form factor at these higher frequencies restricts the advantages of adding more antennas, which diminishes the efficiency gains compared to the upper mid-band. While larger mmWave arrays could in theory achieve even higher gains, such arrays are less feasible in practice due to increased cost and complexity. Hence, gMIMO in the upper mid-band emerges as a promising solution, offering an optimal balance between hardware complexity and enhanced spatial multiplexing capabilities.
@@
Figure 2: Average harvested power versus the number of ID users for different bands.
@@
To illustrate how near-field beamfocusing enhances channel properties and increases EH capability, Fig. 3 compares the far-field and near-field scenarios under perfect CSI assumption. For a fair comparison of the channel structure, all channel vectors are scaled to correspond to a path loss at a distance of $25$ m. Thus, the observed differences in Fig. 3 arise solely from the planar and spherical wave characteristics inherent to the far-field and near-field regions, respectively. All users are positioned within a narrow region near the broadside direction of the gMIMO array to highlight the advantages of near-field massive spatial multiplexing. As shown in Fig. 3, under identical path loss conditions, near-field propagation enables greater average harvested energy. This improvement stems from the more efficient utilization of transmit power and the increased spatial multiplexing capability offered by near-field focusing. With more focused beams, the system can more easily satisfy the SE requirements of ID users using less power, which leaves more power available for EH.
@@
For systems using the power splitting protocol, it is crucial to develop adaptive algorithms that dynamically adjust the power splitting ratio in real time based on battery levels, channel conditions, data requirements, and environmental factors. The SE level can be adaptively adjusted based on users’ remaining energy and predicted traffic patterns. By employing reinforcement learning (RL), the system can anticipate future packet activity and dynamically optimize the power-splitting ratio and precoding strategy. For example, in a smart factory, user activity often follows predictable patterns, which can be effectively learned by RL agents to enable proactive resource adaptation.
@@
The hardware costs of a gMIMO SWIPT system can be analyzed from both the BS and user terminal perspectives. At the BS, the presence of a gigantic number of antennas increases hardware complexity and processing costs compared to conventional massive MIMO systems. Fully digital beamforming is envisioned as the preferred approach to fully exploit the near-field beamfocusing and massive spatial multiplexing capabilities of gMIMO. However, cost-efficient hybrid beamforming architectures, including recently proposed tri-hybrid structures, can also be explored to reduce the number of RF chains while aiming to preserve most of the near-field processing gain. Designing such architectures to balance hardware cost and near-field performance constitutes an important research question for future work. On the user side, ID user terminals equipped with a greater number of antennas also experience increased hardware costs. Similarly, EH users can benefit from multiple rectennas—antennas that rectify signals— which enhance harvested energy but bring additional hardware costs.
@@
The upper mid-band frequencies exhibit sparser propagation characteristics compared to sub-6 GHz systems, which simplifies channel estimation using parametric techniques. In the considered near-field channel model, each user’s LOS channel component can be described by three unknown parameters: one distance and two angular variables. As a result, the number of required pilot symbols is on the same order as the number of unknown parameters—typically only a few symbols per user—rather than scaling with the number of antennas as in conventional LS estimation. This leads to a substantial reduction in pilot overhead for large gMIMO arrays. Classical direction-of-arrival algorithms and compressed sensing-based techniques can be effectively employed to estimate these few channel parameters.
@@
To observe the effect of channel estimation errors on the harvested energy, we consider a gMIMO system with $20$ ID users and $5$ EH users in Fig. 4. Under the requirement for SE of $4$ bit/s/Hz for the ID users, the cumulative distribution function (CDF) of the net average harvested power per EH user is plotted. Initially, the EH users transmit orthogonal pilot sequences of $5$ symbols to the gMIMO BS, which then estimates their channels using either a parametric maximum likelihood (ML)-based channel estimator or a least-squares (LS) channel estimator. The analog circuit power consumption of each EH device is set to $3$ mW, consistent with the low-power IoT device specifications in [15]. As shown in Fig. 4, the CDF curves corresponding to the ML estimator are shifted further to the right, indicating that more power can be harvested when using the ML estimator, as it provides higher channel estimation accuracy. However, this improvement comes at a significant computational cost due to the three-dimensional grid search required in ML estimation, which makes the number of operations scale cubically with the grid resolution. In contrast, the LS estimator involves only a straightforward matrix multiplication. Striking a balance between estimation accuracy and computational complexity remains an important research direction.
@@
In the lower part of Fig. 5, we consider near-field SWIPT with a higher number of ID user antennas, as indicated in the legend. Interestingly, increasing the number of ID user antennas decreases the maximum average harvested power showing a specific trade-off relationship in the near-field multi antenna SWIPT scenarios. This can be attributed to the fact that the energy beamformers are designed in the null space of the ID precoding matrix, which has a higher dimension as the number of subchannels increases in a near-field scenario. Consequently, the dimensionality of the space available for the energy precoder becomes more restricted when the number of antennas at the ID user increases. This observation is noteworthy, as the optimal number of antennas may vary depending on the SE requirement of the robotic arm. For lower SE requirements, fewer antennas can be activated, which leads to a modified version of the classical water-filling algorithm. This adaptation makes optimization for multi-antenna users an intriguing research direction.
@@
Sparse arrays have been extensively studied in the field of array signal processing. In a sparse array, sensors or antennas are arranged in a (non)uniform manner, with part of/all elements spaced more than half a wavelength apart. This design allows for an increased array aperture with a fixed number of antennas, improving the array resolution. Sparse arrays also offer a potential solution to the hardware cost challenges of gMIMO systems, where the goal is to deploy antennas within a given array aperture to preserve the near-field effects for SWIPT capabilities of the gMIMO array. By adopting sparse array designs, the cost can be reduced compared to dense antenna placement within the array aperture. Furthermore, sparse arrays can be dynamically configured by selecting antennas from a dense antenna grid based on real-time CSI and the locations of EH and ID users. In industrial environments, ceiling-mounted sparse arrays may be effectively realized by selecting an appropriate subset of elements from a dense grid, where the choice is guided by real-time CSI and the spatial distribution of EH and ID users, thereby enabling dynamic adaptation to the prevailing network conditions.
As previously mentioned, modular arrays (also referred to as modular MIMO) use groups of densely packed antennas arranged sparsely across the aperture. This structure is particularly suitable for distributed deployments, such as modules mounted along a building facade or on factory pillars. Beyond hardware cost reduction, modular arrays enable the development of novel distributed SWIPT precoding algorithms by leveraging cost-efficient local processing at each subarray. Future research should also explore advanced rectennas with higher RF-to-DC conversion efficiency to further reduce energy consumption while maintaining performance.